A coin is flipped 3 times. If you get no tail you win nothing, tails once you get $5, tails
twice you get $15, and tails on all three flips you get $40.
a. List the probability distribution of the amount you win
b. What is the expected amount of money you will get?

Found 2 solutions by greenestamps, saw:
Answer by greenestamps(13198)



HHH

HHT
HTH
TTH

HTT
THT
TTH

TTT

a. Probabilities....

0 tails: 1 way out of 8 total: P(0) = 1/8
1 tails: 3 ways out of 8 total: P(1) = 3/8
2 tails: 3 ways out of 8 total: P(2) = 3/8
3 tails: 1 way out of 8 total: P(3) = 1/8

b. Expected payout



The expected winnings are $12.50.


Answer by saw(34)

1.HHH
2.HHT
3.HTH
4.HTT
5.TTT
6.TTH
7.THT
8.THH
No tail = 1/8
One tail = 3/8
Two tails = 3/8
All tails = 1/8
(1/8×0)+(3/8×5)+(3/8×15)+(1/8×40) = 12.5